# AraDiCE: Benchmarks for Dialectal and Cultural Capabilities in LLMs

### v1

@@ 6 Related Work
There has been considerable work on benchmarking language models in Arabic, with prior research often including Arabic in multilingual benchmarks like XGLUE Liang et al. (2020), XTREME Hu et al. (2020), XTREME-R Ruder et al. (2021), GEM Gehrmann et al. (2021), Dophin Nagoudi et al. (2023). These benchmarks cover a variety of tasks, primarily focusing on classification problems such as natural language inference and generation, part-of-speech tagging, named entity recognition, and summarization. Another relevant dataset is EXAMS Hardalov et al. (2020), featuring multilingual school exam questions, including about 500 in Arabic. This dataset extends evaluation to educational contexts, providing a diverse testing ground for language models. Notable benchmarks in the realm of question answering (QA), that primarily aim to evaluate models on reading comprehension and QA include TyDiQA Clark et al. (2020), Arabic-SQuAD Mozannar et al. (2019), and MLQA Lewis et al. (2020). These datasets primarily aim to evaluate models on reading comprehension and QA. More recently, Koto et al. (2024) introduced ArabicMMLU, tailored to the specific needs of the MENA region. The dataset includes 40 tasks in MSA. However, a limitation of their work, as they acknowledged, is the lack of focus on Dialectal Arabic. There has been considerable effort in training Arabic language models (LMs) Antoun et al. (2020); Inoue et al. (2021); Abdul-Mageed et al. (2021); Nagoudi et al. (2022) and, more recently, LLMs such as Jais Sengupta et al. (2023) and AceGPT Huang et al. (2024). Jais is pre-trained from scratch using 72 billion Arabic tokens, while AceGPT builds on Llama2 using reinforcement learning from AI feedback. There has also been efforts benchmarking LLMs for Arabic, most notably focusing on standard LLMs based dataset evaluation Sengupta et al. (2023) and NLP tasks Khondaker et al. (2023); Abdelali et al. (2024), and multimodality Alwajih et al. (2024); Das et al. (2024). Focusing on benchmarking the cultural capabilities of LLMs, this includes measuring how entity mentions are culturally biased towards Western or Arab contexts Naous et al. (2024), and assessing cultural alignment based on the World Values Survey AlKhamissi et al. (2024). Other notable work on cultural benchmarking include Demidova et al. (2024); Shen et al. (2024); Liu et al. (2024); Arora et al. (2024); Myung et al. (2024). While prior work has focused on benchmarking LLMs for MSA, our study extends this by evaluating both MSA and dialects (Lev and Egy) using Arabic-centric and non-Arabic-centric LLMs.
@@ 7 Conclusions
We develop benchmarks for dialectal Arabic using machine translation and human post-editing. Our benchmarks evaluate various NLP tasks, including understanding, generation, and cultural awareness across Gulf, Lev, and Egy dialects. Our results show that while Arabic-centric models like Jais and AceGPT perform better in dialectal contexts, they still face challenges compared to MSA and English. Performance varies by dialect and task, highlighting the need for more specialized training for effective handling of regional linguistic and cultural nuances. We release our dialectal benchmarks and models to support future research and advancements in NLP for low-resource languages. We will release the dialectal translation models and benchmarks developed in this study to support further research in NLP for low-resource languages.
@@ A.1 Understanding and Generation
  • Machine Translation: The MADAR corpus Bouamor et al. (2018b) is used for machine translation tasks, providing translations between 25 Arabic city-level dialects and other languages.
@@ B.1 Data
  • Madar Bouamor et al. (2018a): Parallel corpus of 25 Arabic city dialects in the travel domain. It comprises data in the Levantine, Egyptian, Moroccan, and Gulf dialects.
@@
In some cases, the dataset does not contain a dialectal counterpart. To address this issue, we translate the data from English to the corresponding dialect using the NLLB (nllb-200-3.3B) [^14] https://huggingface.co/facebook/nllb-200-3.3B base model Team et al. (2022). The number of parallel sentences included in each dataset is presented in Table 4.
@@ B.3 Machine Translation Results
In our preliminary experiments, we found that the NLLB 3.3B model outperformed AraT5 and its smaller variants after fine-tuning with dialectal data (Table 7). We carried out ablation studies using different data mixtures on the NLLB 3.3B model. Initially, we shortlisted three systems per dialect using BLEU scores as our primary criterion (Tables 5, 6). Nonetheless, basing decisions solely on BLEU scores presents a risk of over-fitting. We observed improved performance on specific tests after incorporating in-domain training data. Therefore, we carried out a human evaluation of the selected systems to select the best model for translation. For this, we selected a sample of 100 sentences from different genres, and carried out translations using the systems. We conducted a human evaluation by showing the output of three translation systems and asked the participants to select the “best system(s)”. [^15] Note that the translations were presented in a Google Doc. The instructions to the annotators were as follows:
@@
In Tables 8 and 9, we report our human evaluation results, which show that System S1, which is NLLB tuned just using UFAL data was best for translating MSA-to-Levantine and System S2 trained using MADAR, Dial2MSA and LDC data were better. However, in our human evaluation, we found that humans had more confidence in the MSA-to-Levantine system than in MSA-to-Egyptian.
@@
Table 8: Human evaluation of different systems across datasets. None refers that the output from none of the systems were acceptable.
@@
Table 9: Human evaluation for the selected MSA-to-EGY models.
@@ C.1 Dialect Identification
The results of the dialect identification task are shown in Table 10. In Figures 10 and 11 we present the confusion matrices for the ADI and ADD datasets, respectively. For ADI, Llama 3, Mistral, and AceGPT tend to confuse MSA with the Levantine and Gulf dialects, while Jais performs best in detecting the Egyptian dialect. For ADD, Llama 3 and Jais excel at detecting the Egyptian dialect, whereas AceGPT confuses all dialects with MSA. The variation in model performance is attributed to the nature of the data and its underlying distribution.
@@ C.3 MT as Generation
Figure 12: ArabicMMLU categories. Arabic alphabets prompt.
Figure 13: Average scores on ArabicMMLU including larger LLM.
@@
Dataset Llama 3 Mistral Jais AceGPT
Gulf 4.35 0.16 6.59 3.62
Levantine 4.22 0.09 6.19 3.32
Egyptian 4.58 0.12 6.56 3.31
Table 19: Averaged results for the dialectal to MSA systems across models.
@@ D.1 World Knowledge
The results in Figure 12 provide key insights into the performance of different models on MSA and dialectal datasets (Egyptian and Levantine). Below we provide some key observations.
@@ F.2.1 ArabicMMLU
Below are a few examples from different disciplines and specific instructions for scenarios. We show the questions and answers (options). We provide the source (MSA), MT, and PEMT Levantine for illustration purposes. However, for Egyptian, the instructions are the same for the dialectal translation. For each example, we provide an explanation highlighting the reason for the edits. 1. Islamic Studies Question (MSA): {RLtext}قال تعالى في سورة العاديات: ( فالمغيرات صبحا (3)) ماالمقصود بالمغيرات؟ Options (MSA): {RLtext}الرياح الشديدة 1. {RLtext}الابل 2. {RLtext}الملائكة السابحين 3. {RLtext}الخيل 4. Question (PEMT -Levantine):{RLtext}قال تعالى في سورة العاديات: ( فالمغيرات صبحا (3)) شو يعني المغيرات؟ Options (PEMT -Levantine): {RLtext}الهوا القوي 1. {RLtext}الابل 2. {RLtext}الملائكة السابحين 3. {RLtext}الخيل 4. Explanation: Verses from Qur’an should not be modified. Any changes to such elements will alter the meaning of the question. However, in this case, the options can be expressed differently in the dialect. 2. Accounting Question (MSA): {RLtext}فرح محاسبي يهدف إلى التحقق من صحة وسلامة المعلومات المالية بما يضفى مزيدًا من الثقة على هذه المعلومات Options (MSA): {RLtext}المحاسبة الحكومية 1. {RLtext}المحاسبة الدولية 2. {RLtext}المراجعة 3. {RLtext}المحاسبة الاجتماعية 4. Question (PEMT -Levantine): {RLtext}فرع محاسبي بيهدف للتحقق من من صحة وسلامة المعلومات المالية وبالتالي بزيد الثقة من هيدي المعلومات Options (PEMT -Levantine): {RLtext}المحاسبة الحكومية 1. {RLtext}المحاسبة الدولية 2. {RLtext}المراجعة 3. {RLtext}المحاسبة الاجتماعية 4. Explanation: The names of knowledge domains and sub fields (the options) should remain the same. Changes should only be made to correct spelling mistakes in these words and phrases. 3. Arabic Language Question (MSA): {RLtext}الكلمة التي تحتوي على حرف السين(س) في آخر الكلمة: Options (MSA): {RLtext}فرس 1. {RLtext}سلمى 2. {RLtext}مسلم 3. Question (PEMT -Levantine): {RLtext}الكلمة اللي بآخرها حرف السين(س): Options (PEMT -Levantine): {RLtext}فرس 1. {RLtext}فرس 2. {RLtext}مسلم 3. Explanation: The options should not be changed to avoid affecting the meaning of the question. 4. Arabic Language (Grammar) Question (MSA): {RLtext}اختر حرف الجر المناسب لوضعه بدلاً من [فراغ] في الجملة التالية: أنا [فراغ] باكستان. Options (MSA): {RLtext}ب 1. {RLtext}إلى 2. {RLtext}من 3. {RLtext}في 4. Question (PEMT -Levantine): {RLtext}نقي حرف الجر المناسب لتحطته مطرح [فراغ] بهيدي الجملة: أنا [فراغ] باكستان. Options(PEMT -Levantine): {RLtext}ب 1. {RLtext}إلى 2. {RLtext}من 3. {RLtext}في 4. 5. Biology Question (MSA): {RLtext}أي من الأت مٌثل الكرموسومات الأنثو ةٌ ؟ Question (PEMT -Levantine): {RLtext}اي احتمال بمثل كرموسومات الأنثوة؟ Options:

### v3

@@ 6 Related Work
There has been considerable work on benchmarking language models in Arabic, with prior research often including Arabic in multilingual benchmarks like XGLUE Liang et al. (2020), XTREME Hu et al. (2020), XTREME-R Ruder et al. (2021), GEM Gehrmann et al. (2021), Dophin Nagoudi et al. (2023). These benchmarks cover a variety of tasks, primarily focusing on classification problems such as natural language inference and generation, part-of-speech tagging, named entity recognition, and summarization. Another relevant dataset is EXAMS Hardalov et al. (2020), featuring multilingual school exam questions, including about 500 in Arabic. This dataset extends evaluation to educational contexts, providing a diverse testing ground for language models. Notable benchmarks in the realm of question answering (QA), that primarily aim to evaluate models on reading comprehension and QA include TyDiQA Clark et al. (2020), Arabic-SQuAD Mozannar et al. (2019), and MLQA Lewis et al. (2020). These datasets primarily aim to evaluate models on reading comprehension and QA. More recently, Koto et al. (2024) introduced ArabicMMLU, focusing on the MENA region. The dataset includes 40 tasks in MSA. However, a limitation of their work, is the lack of dialectal Arabic. Significant efforts have been made in training Arabic language models (LMs) Antoun et al. (2020); Inoue et al. (2021); Abdul-Mageed et al. (2021); Nagoudi et al. (2022) and more recently, LLMs like Jais Sengupta et al. (2023), AceGPT Huang et al. (2024), ALLaM Bari et al. (2024), Arabic Stable LM Alyafeai et al. (2024), and FANAR Fanar-Team (2024). Jais is trained from scratch on 72 billion Arabic tokens, AceGPT builds on Llama2 with reinforcement learning from AI feedback, ALLaM spans 7B–70B parameters but is not publicly available, Arabic Stable LM fine-tunes Stable LM 2 with 1.6B parameters, and FANAR LLM supports diverse language, speech, and image generation tasks. There has also been efforts benchmarking LLMs for Arabic, most notably focusing on standard LLMs based dataset evaluation Sengupta et al. (2023) and NLP tasks Khondaker et al. (2023); Abdelali et al. (2024); Dalvi et al. (2024), prompting in native (Arabic) and non-native (English) language Kmainasi et al. (2025) and multimodality Alwajih et al. (2024); Das et al. (2024). Focusing on benchmarking the cultural capabilities of LLMs, this includes measuring how entity mentions are culturally biased towards Western or Arab contexts Naous et al. (2024), and assessing cultural alignment based on the World Values Survey AlKhamissi et al. (2024). Other notable work on cultural benchmarking include Demidova et al. (2024); Shen et al. (2024); Liu et al. (2024); Arora et al. (2024); Myung et al. (2024). While prior work has focused on benchmarking LLMs for MSA, our study extends this by evaluating both MSA and dialects (Lev and Egy) using Arabic-centric and non-Arabic-centric LLMs.
@@ 7 Conclusions
We developed dialectal Arabic benchmarks through machine translation and post-editing. Our benchmarks evaluate various NLP tasks, including understanding, generation, and cultural awareness across Arabic dialects. Our results show that while Arabic-centric models like FANAR, Jais and AceGPT perform better in dialectal contexts, they still face challenges compared to MSA and English. Performance varies by dialect and task, highlighting the need for more specialized training for effective handling of regional linguistic and cultural nuances. We have released our dialectal benchmarks and models to support future research and advancements in NLP for low-resource languages. We also released the dialectal translation models [^8] https://huggingface.co/datasets/QCRI/AraDiCE and benchmarks developed in this study to support further research in NLP for low-resource languages.
@@ A.1 Understanding and Generation
  • Machine Translation: The MADAR corpus Bouamor et al. (2018) is used for machine translation tasks, providing translations between 25 Arabic city-level dialects and other languages.
@@ Appendix B Machine Translation
Machine translation of Arabic is challenging due to morphological complexity and dialectal variations Birch et al. (2014); Sajjad et al. (2017). Here we detail our efforts in training MSA–dialect models.
@@ B.1 Data
  • Madar Bouamor et al. (2018): Parallel corpus of 25 Arabic city dialects in the travel domain. It comprises data in the Levantine, Egyptian, Moroccan, and Gulf dialects.
@@
In some cases, the dataset does not contain a dialectal counterpart. To address this issue, we translate the data from English to the corresponding dialect using the NLLB (nllb-200-3.3B) [^14] https://huggingface.co/facebook/nllb-200-3.3B base model Team et al. (2022). The number of parallel sentences included in each dataset is presented in Table 4. We used dialectal tests in AraBench Sajjad et al. (2020) for our evaluation.
@@ B.3 Machine Translation Results
In our preliminary experiments, we found that the NLLB 3.3B model outperformed AraT5 and its smaller variants after fine-tuning with dialectal data (Table 7). We carried out ablation studies using different data mixtures on the NLLB 3.3B model. Initially, we shortlisted three systems per dialect using BLEU scores as our primary criterion (Tables 5, 6). Nonetheless, basing decisions solely on BLEU scores presents a risk of over-fitting. We observed improved performance on specific tests after incorporating in-domain training data. Therefore, we carried out a human evaluation of the selected systems to select the best model for translation. For this, we selected a sample of 100 sentences from different genres, and carried out translations using the systems. We conducted a human evaluation by showing the output of three translation systems and asked the participants to select the “best system(s)”. The instructions to the annotators were as follows:
@@
In Tables 8 and 9, we report our human evaluation results, which show that System S1, which is NLLB tuned just using UFAL data was best for translating MSA-to-Levantine and System S2 trained using MADAR, Dial2MSA and LDC data was better for translating MSA-to-Egyptian translation. However, in our human evaluation, we found that humans had more confidence in the MSA-to-Levantine system than in MSA-to-Egyptian.
@@
Table 8: Human evaluation of different MSA-to-LEV systems across datasets. None refers that the output from none of the systems were acceptable.
@@
Table 9: Human evaluation for the selected MSA-to-EGY models.None refers that the output from none of the systems were acceptable.
@@ C.1 Dialect Identification
The results of the dialect identification task are shown in Table 10. In Figures 12 and 13 we present the confusion matrices for the ADI and ADD datasets, respectively. For ADI, Llama 3, Mistral, and AceGPT tend to confuse MSA with the Levantine and Gulf dialects, while Jais performs best in detecting the Egyptian dialect. For ADD, Llama 3 and Jais excel at detecting the Egyptian dialect, whereas AceGPT confuses all dialects with MSA. The variation in model performance is attributed to the nature of the data and its underlying distribution.
@@ C.3 MT as Generation
Figure 14: Results on different subject areas in ArabicMMLU.
@@ D.1 World Knowledge
The results in 19 The results show the overall Arabic MMLU accuracies. in Figure 14 provide key insights into the performance of different models on MSA and dialectal datasets (Egyptian and Levantine). Below we provide some key observations.
@@
MSA LEV EGY
Rand 0.26 0.20 0.20
LLama 3 0.49 0.44 0.44
Mistral 0.40 0.34 0.35
Jais 0.56 0.52 0.51
Ace-GPT 0.57 0.51 0.52
Table 19: Results on ArabicMMLU.
Figure 15: Average results on ArabicMMLU
D.2 Reading Comprehension
The Accuracy numbers for BoolQ and Belebele are shown in tables 20 and 21
MSA LEV EGY ENG
Random 0.38 0.38 0.38 0.38
LLama 3 0.74 0.71 0.73 0.85
Mistral 0.66 0.63 0.63 0.85
Jais 0.71 0.65 0.64 0.75
AceGPT 0.77 0.68 0.69 0.82
Table 20: Results on BoolQ.
MSA LEV EGY ENG
Random 0.26 0.26 0.26 0.23
Llama 3 0.49 0.42 0.39 0.87
Mistral 0.41 0.36 0.35 0.75
Jais 0.57 0.49 0.49 0.80
AceGPT 0.49 0.43 0.42 0.83
Table 21: Results on Belebele.
D.3 Commonsense Reasoning
The results for PiQA, Openbooks, and Winogrande are shown in tables 22, 23, and 24
MSA LEV EGY ENG
Random 0.48 0.49 0.49 0.50
Llama 3 0.55 0.54 0.55 0.78
Mistral 0.55 0.52 0.54 0.81
Jais 0.66 0.59 0.62 0.77
AceGPT 0.64 0.57 0.61 0.80
Table 22: Results on PIQA.
MSA LEV EGY ENG
Random 0.28 0.26 0.31 0.27
Llama 3 0.32 0.33 0.32 0.42
Mistral 0.31 0.30 0.30 0.45
Jais 0.35 0.34 0.32 0.41
AceGPT 0.36 0.34 0.32 0.45
Table 23: Results on OBQA.
MSA LEV EGY
Random 0.51 0.51 0.49
Llama 3 0.60 0.52 0.56
Mistral 0.57 0.51 0.55
Jais 0.65 0.58 0.61
AceGPT 0.61 0.58 0.63
Table 24: Results on Winogrande.
D.4 Misinformation
The results for truthfulQA are shown in table 25
MSA LEV EGY ENG
Random 0.21 0.23 0.20 0.39
Llama 3 0.34 0.29 0.33 0.40
Mistral 0.33 0.33 0.30 0.57
Jais 0.28 0.28 0.31 0.28
AceGPT 0.34 0.30 0.33 0.37
Table 25: Results on TruthfulQA.
@@ F.2.1 ArabicMMLU
Below are a few examples from different disciplines and specific instructions for scenarios. We show the questions and answers (options). We provide the source (MSA), MT, and PEMT Levantine for illustration purposes. However, for Egyptian, the instructions are the same for the dialectal translation. For each example, we provide an explanation highlighting the reason for the edits. 1. Islamic Studies Question (MSA): {RLtext}قال تعالى في سورة العاديات: ( فالمغيرات صبحا (3)) ماالمقصود بالمغيرات؟ Options (MSA): {RLtext}الرياح الشديدة 1. {RLtext}الابل 2. {RLtext}الملائكة السابحين 3. {RLtext}الخيل 4. Question (PEMT -Levantine):{RLtext}قال تعالى في سورة العاديات: ( فالمغيرات صبحا (3)) شو يعني المغيرات؟ Options (PEMT -Levantine): {RLtext}الهوا القوي 1. {RLtext}الابل 2. {RLtext}الملائكة السابحين 3. {RLtext}الخيل 4. Explanation: Verses from Qur’an should not be modified. Any changes to such elements will alter the meaning of the question. However, in this case, the options can be expressed differently in the dialect. 2. Accounting Question (MSA): {RLtext}فرع محاسبي يهدف إلى التحقق من صحة وسلامة المعلومات المالية بما يضفى مزيدًا من الثقة على هذه المعلومات Options (MSA): {RLtext}المحاسبة الحكومية 1. {RLtext}المحاسبة الدولية 2. {RLtext}المراجعة 3. {RLtext}المحاسبة الاجتماعية 4. Question (PEMT -Levantine): {RLtext}فرع محاسبي بيهدف للتحقق من من صحة وسلامة المعلومات المالية وبالتالي بزيد الثقة من هيدي المعلومات Options (PEMT -Levantine): {RLtext}المحاسبة الحكومية 1. {RLtext}المحاسبة الدولية 2. {RLtext}المراجعة 3. {RLtext}المحاسبة الاجتماعية 4. Explanation: The names of knowledge domains and sub fields (the options) should remain the same. Changes should only be made to correct spelling mistakes in these words and phrases. 3. Arabic Language Question (MSA): {RLtext}الكلمة التي تحتوي على حرف السين(س) في آخر الكلمة: Options (MSA): {RLtext}فرس 1. {RLtext}سلمى 2. {RLtext}مسلم 3. Question (PEMT -Levantine): {RLtext}الكلمة اللي بآخرها حرف السين(س): Options (PEMT -Levantine): {RLtext}فرس 1. {RLtext}فرس 2. {RLtext}مسلم 3. Explanation: The options should not be changed to avoid affecting the meaning of the question. 4. Arabic Language (Grammar) Question (MSA): {RLtext}اختر حرف الجر المناسب لوضعه بدلاً من [فراغ] في الجملة التالية: أنا [فراغ] باكستان. Options (MSA): {RLtext}ب 1. {RLtext}إلى 2. {RLtext}من 3. {RLtext}في 4. Question (PEMT -Levantine): {RLtext}نقي حرف الجر المناسب لتحطته مطرح [فراغ] بهيدي الجملة: أنا [فراغ] باكستان. Options(PEMT -Levantine): {RLtext}ب 1. {RLtext}إلى 2. {RLtext}من 3. {RLtext}في 4. 5. Biology Question (MSA): {RLtext}أي من الأت مٌثل الكرموسومات الأنثو ةٌ ؟ Question (PEMT -Levantine): {RLtext}اي احتمال بمثل كرموسومات الأنثوة؟ Options: